OpenStudy (anonymous):

the perimeter of a rectangular garden is 340 feet. if the width of the garden is 76 feet. what is its length??

OpenStudy (yttrium):

It's easy. Take a look at the formula. P = 2L + 2W Given are: P = 340 W = 76 Just plug in the values and solve for L. :)

OpenStudy (shamil98):

340 = 2(76) + 2L solve for L

OpenStudy (anonymous):

L=94

OpenStudy (anonymous):

94... there are two sides eaqual too 76 feet. so 76*2=152. 340-152=188.. you have two remaining sides so 188/2=94

OpenStudy (yttrium):

@ashleynicolee you actually got it right. :))

OpenStudy (shamil98):

Yeah L is 94, good job.